I am having a difficult time understanding this problem. Can someone please help?
A researcher is interested in estimating the proportion of voters who favor a tax on e-commerce. In a sample of 250 people she finds that 36 favor such a tax.
a. Find a 95% confidence interval for the true population proportion of voters who favor such a tax.
b. Find a 99% confidence interval for the true population proportion of voters who favor such a tax.

Answer by stanbon(75887)

I am having a difficult time understanding this problem. Can someone please help?
A researcher is interested in estimating the proportion of voters who favor a tax on e-commerce. In a sample of 250 people she finds that 36 favor such a tax.
---
In favor of:
sample proportion = phat = 36/250 = 0.144
standard error = invNorm(0.975)*sqrt[0.144*0.856/250] = 0.0435..
--------------------
a. Find a 95% confidence interval for the true population proportion of voters who favor such a tax.
95% CI : 0.144-0.0435 < p < 0.144 + 0.0435
--------------------------------------------
b. Find a 99% confidence interval for the true population proportion of voters who favor such a tax.
sample proportion = 0.144
standard error = invNorm(0.995)*sqrt(0.144*0.856/250] = 0.0572
99% CI : 0.144 - 0.0572 < p < 0.144 + 0.0572
==================================================
Cheers,
Stan H.